Mr. Lee is 27 years older than his son. In ten years,he will be twice as old as his son how old each?
Answer by ptaylor(1894)

Let x=son's present age; x+10=son's age in 10 years
Then x+27=Mr. Lee's present age; x+37=Mr. Lee's age in 10 years
Now we are told the following:
x+37=2(x+10) get rid of parens
x+37=2x+20 subtract 20 and also x from each side
x-x+37-20=2x-x+20-20 collect like terms
17=x or
x=17---------------------son's present age
x+27=17+27=44----------------Mr. Lee's present age
CK
44-17=27
27=27
also
44+10=2*27
54=54
Hope this helps----ptaylor